OpenStudy (anonymous):

hi, have couple integral problems need to be answered using U-substitution, can someone help me?

OpenStudy (anonymous):

1) \[\int\limits_{}^{}\frac{ 2x-1 }{ \sqrt{3x ^{2}-3x} } dx\]

OpenStudy (anonymous):

i know the answer of this one but i dont know the steps to answer it correctly

OpenStudy (anonymous):



OpenStudy (anonymous):

Notice that derivative of \(3x^2-3x=6x-3=3(2x-1)\) so you substitution could be \(u=3x^2-3x\) and \(du=3(2x-1)dx\) So the integral will become: \(1/3\int\huge \frac{du}{\sqrt{u}}\)

OpenStudy (anonymous):

so is my answer correct? this is what i came up with after getting the antiderivative and substitution :\[\frac{ 2 }{ 3 } \sqrt{3x ^{2} - 3x} + c\]

OpenStudy (anonymous):

Yes you are right

OpenStudy (anonymous):

Thanks @myko , how about this one: \[\int\limits_{}^{} x^{3} \sqrt{x ^{2}+1}\] \[u= x ^{2} +1 ?\]

OpenStudy (anonymous):

you could go like this: \(u= x^2\), and \(du=2xdx\) so integral becomes: \(\int u\sqrt{u+1}du\) now make one more substitution: \(v=u+1\) and \(dv=du\) integral becomes: \(\int (v-1)\sqrt{v}dv=\int (v^{3/2}-v^{1/2})dv\) which is straight forward

OpenStudy (anonymous):

sry forgot 1/2 infront of integral after 1º subsitution. Because du=2xdx

OpenStudy (anonymous):

i wrote a long solution then the browser crashed :( anyway, did the problem and came out with this answer \[\frac{ 1 }{ 5 } ( x ^{2}+1)^{\frac{ 5 }{ 2 }} - \frac{ 2 }{ 3 }(x ^{2}+1)^{\frac{ 3 }{ 2 }} + c\]

OpenStudy (anonymous):

i guess , to lasy to check :). Looks similar, just that 2/3 i tink should be 1/3

OpenStudy (anonymous):

how about this one: \[\int\limits_{1}^{9} (\sqrt{x}-2)dx\]

OpenStudy (anonymous):

what's the problema with this. \(\int_1^9x^{1/2}dx-\int_1^92dx\)

OpenStudy (anonymous):

what "u" will be?

OpenStudy (anonymous):

you dont need it

OpenStudy (anonymous):

4/3?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

thank you